So far we have had problems with the laser
We also had (2011-06-01) what appeared to be a problem with the scan mirrors / controllers but was actually a software configuration issue.
Testing out the piezo objective positioner loan from PI, we noticed a severe resonance problem (2011-06-03); this was evident in both closed and open loop, but significantly worse in closed loop. Part of the problem appeared to result from having the piezo controller cables running next to the PMT cables. Rotating the piezo controller so that it was sticking out to the left hand side of the objective, rather than underneath the MOM scan head seemed to help.
NI boards could not be accessed from PC (2019-05-17)
In the first case uninstalling the 220A Software, reinstalling NI drivers (version 9.3) solved the problem (the 220A software was reinstalled to a different location). In the second case reinstalling the drivers didn't work, but noticed that the NI Device Loader is not started automatically (check at services.msc). The suggested solution can be found here: https://forums.ni.com/t5/Measure/Service-NI-Device-Loader-isn-t-starting-automatically/td-p/551763?profile.language=en ; https://knowledge.ni.com/KnowledgeArticleDetails?id=kA00Z0000019YidSAE&l=en-GB
This problem couldn't be resolved, updated the PC and the install versions. See 2P1 log - 2019. 09. 09.
With new install (Scanimage 5.6):